# A Linearly Convergent Frank-Wolfe-type Method for Smooth Convex Minimization over the Spectrahedron

### v1

@@ Appendix A Proof of Lemma 2
Let ${\mathbf{X}}\in\mathbb{S}^{n}_{+}$ and ${\mathbf{v}}\in\textrm{Im}({\mathbf{X}})$. Consider the matrix ${\mathbf{Y}}={\mathbf{X}}-\lambda{\mathbf{v}}{\mathbf{v}}^{\top}$ for some $\lambda\geq 0$. If $\lambda\leq({\mathbf{v}}^{\top}{\mathbf{X}}^{\dagger}{\mathbf{v}})^{-1}$, then ${\mathbf{Y}}\succeq 0$. Moreover, if $\lambda=({\mathbf{v}}^{\top}{\mathbf{X}}^{\dagger}{\mathbf{v}})^{-1}$, then $\textrm{rank}({\mathbf{Y}})<\textrm{rank}({\mathbf{X}})$ and ${\mathbf{X}}^{\dagger}{\mathbf{v}}\in\textrm{Ker}({\mathbf{Y}})$.
@@
${\mathbf{Y}}{\mathbf{u}}=\left({{\mathbf{X}}-\frac{1}{{\mathbf{v}}^{\top}{\mathbf{X}}^{\dagger}{\mathbf{v}}}{\mathbf{v}}{\mathbf{v}}^{\top}}\right){\mathbf{X}}^{\dagger}{\mathbf{v}}={\mathbf{X}}{\mathbf{X}}^{\dagger}{\mathbf{v}}-{\mathbf{v}}=0,$
implying that ${\mathbf{u}}\notin\textrm{Im}({\mathbf{Y}})$ and so, $\textrm{rank}({\mathbf{Y}})<\textrm{rank}({\mathbf{X}})$. ∎
@@ Appendix B Fast Update of the Pseudo Inverse Matrix
${\mathbf{B}}^{\dagger}=\left\{\begin{array}[]{ll}{\mathbf{A}}^{\dagger}-{\mathbf{k}}{\mathbf{u}}^{\dagger}-{\mathbf{v}}^{\dagger}{\mathbf{h}}+\zeta{\mathbf{v}}^{\dagger}{\mathbf{u}}^{\dagger}&\mbox{if ${\mathbf{c}}\notin\textrm{Im}({\mathbf{A}})$};\\ {\mathbf{A}}^{\dagger}+\frac{1}{\zeta}{\mathbf{v}}^{\top}{\mathbf{k}}^{\top}{\mathbf{A}}^{\dagger}-\frac{\zeta}{\sigma}{\mathbf{p}}{\mathbf{q}}^{\top}&\mbox{if ${\mathbf{c}}\in\textrm{Im}({\mathbf{A}})\textrm{ and }\zeta\neq 0$};\\ {\mathbf{A}}^{\dagger}-{\mathbf{k}}{\mathbf{k}}^{\dagger}{\mathbf{A}}^{\dagger}-{\mathbf{A}}^{\dagger}{\mathbf{h}}^{\dagger}{\mathbf{h}}+({\mathbf{k}}^{\dagger}{\mathbf{A}}^{\dagger}{\mathbf{h}}^{\dagger}){\mathbf{k}}{\mathbf{h}}&\mbox{if ${\mathbf{c}}\in\textrm{Im}({\mathbf{A}})\textrm{ and }\zeta=0$}.\end{array}\right.$
In particular, given ${\mathbf{A}}^{\dagger}$, one can compute ${\mathbf{B}}^{\dagger}$ in $O(n^{2})$ time.

### v2

@@ 5 Future Directions
6 Acknowledgments
This work was funded by the European Union (ERC, ProFreeOpt, 101170791). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Proof of Lemma 2
Let ${\mathbf{X}}\in\mathbb{S}^{n}_{+}$ and ${\mathbf{v}}\in\textrm{Im}({\mathbf{X}})$, ${\mathbf{v}}\neq\mathbf{0}$. Consider the matrix ${\mathbf{Y}}={\mathbf{X}}-\lambda{\mathbf{v}}{\mathbf{v}}^{\top}$ for some $\lambda\geq 0$. If $\lambda\leq({\mathbf{v}}^{\top}{\mathbf{X}}^{\dagger}{\mathbf{v}})^{-1}$, then ${\mathbf{Y}}\succeq 0$. Moreover, if $\lambda=({\mathbf{v}}^{\top}{\mathbf{X}}^{\dagger}{\mathbf{v}})^{-1}$, then $\textrm{rank}({\mathbf{Y}})=\textrm{rank}({\mathbf{X}})-1$ and ${\mathbf{X}}^{\dagger}{\mathbf{v}}\in\textrm{Ker}({\mathbf{Y}})$.
@@
${\mathbf{Y}}{\mathbf{u}}=\left({{\mathbf{X}}-\frac{1}{{\mathbf{v}}^{\top}{\mathbf{X}}^{\dagger}{\mathbf{v}}}{\mathbf{v}}{\mathbf{v}}^{\top}}\right){\mathbf{X}}^{\dagger}{\mathbf{v}}={\mathbf{X}}{\mathbf{X}}^{\dagger}{\mathbf{v}}-{\mathbf{v}}=\mathbf{0},$
implying that ${\mathbf{u}}\notin\textrm{Im}({\mathbf{Y}})$ and so, $\textrm{rank}({\mathbf{Y}})=\textrm{rank}({\mathbf{X}})-1$. ∎
@@ Appendix B Fast Update of the Pseudoinverse Matrix
$({\mathbf{A}}+{\mathbf{c}}{\mathbf{d}}^{\top})^{\dagger}=\left\{\begin{array}[]{ll}{\mathbf{A}}^{\dagger}-{\mathbf{k}}{\mathbf{u}}^{\dagger}-{\mathbf{v}}^{\dagger}{\mathbf{h}}+\zeta{\mathbf{v}}^{\dagger}{\mathbf{u}}^{\dagger}&\mbox{if ${\mathbf{c}}\notin\textrm{Im}({\mathbf{A}})$};\\ {\mathbf{A}}^{\dagger}+\frac{1}{\zeta}{\mathbf{v}}^{\top}{\mathbf{k}}^{\top}{\mathbf{A}}^{\dagger}-\frac{\zeta}{\sigma}{\mathbf{p}}{\mathbf{q}}^{\top}&\mbox{if ${\mathbf{c}}\in\textrm{Im}({\mathbf{A}})\textrm{ and }\zeta\neq 0$};\\ {\mathbf{A}}^{\dagger}-{\mathbf{k}}{\mathbf{k}}^{\dagger}{\mathbf{A}}^{\dagger}-{\mathbf{A}}^{\dagger}{\mathbf{h}}^{\dagger}{\mathbf{h}}+({\mathbf{k}}^{\dagger}{\mathbf{A}}^{\dagger}{\mathbf{h}}^{\dagger}){\mathbf{k}}{\mathbf{h}}&\mbox{if ${\mathbf{c}}\in\textrm{Im}({\mathbf{A}})\textrm{ and }\zeta=0$}.\end{array}\right$.
In particular, given $\mathbf{A}$$^{\dagger}$, one can compute $({\mathbf{A}}+{\mathbf{c}}{\mathbf{d}}^{\top})^{\dagger}$ in $O(n^{2})$ time.
Appendix C Auxiliary Lemmas
The following lemma is well-known, but for clarity of presentation we include it and the short proof.
Lemma 10.
Let $\Psi:\mathbb{E}\rightarrow\mathbb{R}$ be $\beta_{\Psi}$-smooth and convex over $\mathcal{K}$ — a convex and compact subset of a Euclidean space $\mathbb{E}$. The gradient $\nabla{}\Psi$ is constant over the set of minimizers $\arg\min_{{\mathbf{z}}\in\mathcal{K}}\Psi({\mathbf{z}})$, and for any ${\mathbf{z}}\in\mathcal{K}$ it holds that
$\|{\nabla\Psi({\mathbf{z}})-\nabla\Psi({\mathbf{z}}^{*})}\|^{2}\leq\beta_{\Psi}\left({\Psi({\mathbf{z}})-\Psi({\mathbf{z}}^{*})}\right),$ (34)
where $\mathbf{z}$$^{*}$ is any point in $\arg\min_{{\mathbf{z}}\in\mathcal{K}}\Psi({\mathbf{z}})$.
Proof.
Since $\Psi(\cdot)$ is smooth we have that,
$\|{\nabla\Psi({\mathbf{z}})-\nabla\Psi({\mathbf{z}}^{*})}\|^{2}$ $\leq\beta_{\Psi}\langle{{\mathbf{z}}-{\mathbf{z}}^{*},\nabla\Psi({\mathbf{z}})-\nabla\Psi({\mathbf{z}}^{*})}\rangle$
$\leq\beta_{\Psi}\left({\Psi({\mathbf{z}})-\Psi({\mathbf{z}}^{*})}\right)+\beta_{\Psi}\langle{{\mathbf{z}}^{*}-{\mathbf{z}},\nabla\Psi({\mathbf{z}}^{*})}\rangle$
$\leq\beta_{\Psi}\left({\Psi({\mathbf{z}})-\Psi({\mathbf{z}}^{*})}\right),$
where the second inequality is due to the convexity of $\Psi$, and the last one is due to the first-order optimality condition.
This proves both parts of the lemma. ∎
Lemma 11.
If Assumption 2 holds with $r^{*}\geq 2$, then the smoothness and quadratic growth constants satisfy $\alpha\leq\beta$.
Proof.
Let $\mathbf{X}$$^{*}$ be an optimal solution. Under Assumption 2, we have that $\textrm{rank}({\mathbf{X}}^{*})=r^{*}$, and let us write its eigen-decomposition as ${\mathbf{X}}^{*}={\mathbf{U}}\Lambda{\mathbf{U}}^{\top}$, where ${\mathbf{U}}\in\mathbb{R}^{n\times{}r^{*}}$ and $\Lambda\in\mathbb{R}^{r^{*}\times r^{*}}$ is diagonal and positive definite. Let $\Lambda^{\prime}\in\mathbb{R}^{r^{*}\times{}r^{*}}$ be a diagonal positive semidefinite matrix with unit trace such that $\textrm{rank}(\Lambda^{\prime})<\textrm{rank}(\Lambda)$, and consider the matrix ${\mathbf{X}}={\mathbf{U}}\Lambda^{\prime}{\mathbf{U}}^{\top}\in\mathcal{S}^{n}$. From the first-order optimality condition, see for instance Lemma 5.2 in [14], it follows that $\langle{{\mathbf{X}}-{\mathbf{X}}^{*},\nabla{}f({\mathbf{X}}^{*})}\rangle=0$. Thus, using the smoothness and quadratic growth of $f$ we have that,
$f({\mathbf{X}})$ $\leq f({\mathbf{X}}^{*})+\langle{{\mathbf{X}}-{\mathbf{X}}^{*},\nabla{}f({\mathbf{X}}^{*})}\rangle+\frac{\beta}{2}\|{{\mathbf{X}}-{\mathbf{X}}^{*}}\|_{F}^{2}$
$=f({\mathbf{X}}^{*})+\frac{\beta}{2}\|{{\mathbf{X}}-{\mathbf{X}}^{*}}\|_{F}^{2}$
$\leq f({\mathbf{X}}^{*})+\frac{\beta}{\alpha}\left({f({\mathbf{X}})-f({\mathbf{X}}^{*})}\right).$
Since $\textrm{rank}(\Lambda^{\prime})<\textrm{rank}(\Lambda)=\textrm{rank}({\mathbf{X}}^{*})$, we have that, under Assumption 2, $\mathbf{X}$} is not an optimal solution, meaning $f({\mathbf{X}})>f({\mathbf{X}}^{*})$, which completes the proof. ∎